Please help me solve this equation. x^2-2
------- for x= -1 and x= 3
x^2-9

Answer by checkley77(12844)

x^2-2
------- for x= -1 and x= 3
x^2-9
--------------------------------
f(-1)=(1^2-2)/(1^2-9)
f(-1)=(1-2)/(1-9)
f(-1)=-1/-8 =1/8 answer.
------------------------------
f(3)=(3^2-2)/(3^2-9)
f(3)=(9-2)/(9-9)
f(3)=7/0=(infinity) answer.